OpenStudy (anonymous):

tia's car needs repairs. honest harry will charge $70 per hour plus $130 for the part. lucky lou will charge $80 per hour plus $40 for the part. how long is the job if the two costs are the same?

OpenStudy (anonymous):

70 X + 130 = 80 X + 40 130 - 40 = 80 X - 70 X 90 = 10 X X = 9

OpenStudy (mathstudent55):

Let the number of hours = x Honest Harry: chages 70x + 130 Lucky Lou charges 80x + 40 Set the two charges equal to each other and solve for x.